# The Promises and Pitfalls of LLM Annotations in Dataset Labeling: a Case Study on Media Bias Detection

### v1

@@ Abstract
Tomáš Horych1, Christoph Mandl1, Terry Ruas2, André Greiner-Petter1, Bela Gipp2, Akiko Aizawa1, Timo Spinde2 1National Institute of Informatics, Tokyo, Japan 2University of Göttingen, Göttingen, Germany t.horych@media-bias-research.org, c.mandl@media-bias-research.org, ruas@uni-goettingen.de, greinerpetter@gipplab.org, gipp@uni-goettingen.de aizawa@nii.ac.jp, t.spinde@media-bias-research.org
@@ 5.1 The annotation corpus
Extract. We start with scraping all public articles from outlets that have ratings from both allsides and adfontesmedia platforms. Both platforms use left-to-right ratings with different scales; we unify their ratings into five labels: Left, Lean Left, Center, Lean Right, and Right. We only keep articles where both platforms agree on the rating. Filter. We filter out empty, short, or other corrupted articles and keep only articles written in English. We then segment these articles into sentences. Additionally, we trim special characters and other irregularities from the sentences. The final collection of filtered sentences contains approximately 400,000 sentences. Sample. Finally, We sample sentences to ensure the balance across the aforementioned political spectrum. However, we can’t ensure a fair distribution of lexical bias before knowing the true labels (i.e., before annotation). Some outlets may be more likely to contain lexical bias, which could result in an uneven distribution, with one side of the spectrum having mostly biased sentences and the other side being largely neutral after the annotation.
@@ Acknowledgements
To keep anonymity, acknowledgments will be added in case of acceptance.
@@ Appendix A Prompting
You are an expert in media bias. { for TEXT, LABEL, EXPLANATION in examples } Instruction: ’[TEXT]’ Classify the sentence above as BIASED or NOT BIASED. Output: Let’s think step by step. [EXPLANATION] The answer is [LABEL]. { endfor } Instruction: ’[SENTENCE]’ Classify the sentence above as BIASED or NOT BIASED. Output: Let’s think step by step.

### v2

@@ Abstract
Tomáš Horych1, Christoph Mandl1, Terry Ruas1, André Greiner-Petter2, Bela Gipp1, Akiko Aizawa2, Timo Spinde1 1University of Göttingen, Göttingen, Germany 2National Institute of Informatics, Tokyo, Japan t.horych@media-bias-research.org, c.mandl@media-bias-research.org, ruas@uni-goettingen.de, greinerpetter@gipplab.org, gipp@uni-goettingen.de aizawa@nii.ac.jp, t.spinde@media-bias-research.org
@@ 5.1 The annotation corpus
Extract. We start with scraping all public articles from outlets that have ratings from both allsides and adfontesmedia platforms. Both platforms use left-to-right ratings with different scales; we unify their ratings into five labels: Left, Lean Left, Center, Lean Right, and Right. We only keep articles where both platforms agree on the rating.
Filter. We filter out empty, short, or other corrupted articles and keep only articles written in English. We then segment these articles into sentences. Additionally, we trim special characters and other irregularities from the sentences. The final collection of filtered sentences contains approximately 400,000 sentences.
Sample. Finally, We sample sentences to ensure the balance across the aforementioned political spectrum. However, we can’t ensure a fair distribution of lexical bias before knowing the true labels (i.e., before annotation). Some outlets may be more likely to contain lexical bias, which could result in an uneven distribution, with one side of the spectrum having mostly biased sentences and the other side being largely neutral after the annotation.
@@ Acknowledgements
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Furthermore, this project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 554559555 and some of the results partially by EXIST-Gründungsstipendium. Finally, the authors would like to express their gratitude towards Prof. Dr. Michael Granitzer for consulting and valuable advices.
@@ Appendix A Prompting
You are an expert in media bias.
{ for TEXT, LABEL, EXPLANATION in examples }
Instruction: ’[TEXT]’
Classify the sentence above as BIASED or NOT BIASED.
Output: Let’s think step by step. [EXPLANATION] The answer is [LABEL].
{ endfor }
Instruction: ’[SENTENCE]’
Classify the sentence above as BIASED or NOT BIASED.
Output: Let’s think step by step.